# Price of Uncertainty for Consensus Games

### v1

@@ 3.2 Proof of Improved Upper Bound
With our construction, Lemma 7 allows us to contain the bad edges in the graph inside one of $S_{0},S_{1},\dots,S_{m}$. We will now present a combinatorial abstraction that will allow us to focus on the effect of the first uncertain best response played by vertices in $V$.
We can use $S_{k}$ to construct an infinite sequence $E_{k}$ of positive integers, finitely many of which are nonzero. Each nonzero element of $E_{k}$ will encode the number of edges incident to a vertex in $G\setminus S_{k}$ and any vertex in $S_{k}$, and each vertex in $G\setminus S_{k}$ will be encoded exactly once in such a manner. The rest of $E_{k}$ will be populated with zeros. Finally, to ensure that $E_{k}$ is unique, we will also include a requirement that $E_{k}$ must be sorted in descending order. For example, suppose that there existed 4 edges incident to a vertex in $S_{0}$ and some vertex $x_{1}$, and 3 edges incident to a vertex in $S_{0}$ and some vertex $x_{2}\neq x_{1}$. Also, suppose that there existed no other edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$. Then, we would have
$E_{0}=\{4,3,0,0,0,\dots\}.$
Consider the state of the graph right before $v_{k}$ plays its first uncertain best response. By Lemma 7, all edges incident to $v_{k}$ and a vertex in $G\setminus S_{k-1}$ must be good. If we let $z$ be the number of edges incident to $v_{k}$ and a vertex in $S_{k-1}$, then the degree of $v_{k}$ cannot be greater than $z+\lfloor(1+\varepsilon)z\rfloor$. From this logic, going from $E_{k-1}$ to $E_{k}$ consists of deleting the element of $E_{k-1}$ corresponding to edges incident to a vertex in $S_{k-1}$ and $v_{k}$, which is $z$. We will now select $\lfloor(1+\varepsilon)z\rfloor$ elements of $E_{k-1}$ and increment them by $1$. This represents the new outgoing edges from $v_{k}$ to $G\setminus S_{k}$. Since any graph and sequence of uncertain best responses can be represented using the $E_{k}$ in such a manner, it suffices to study the behavior of these moves.
@@
A move on a sequence $E_{k}$ consists of the following steps:
  • Select some integer $z$ from $E_{k}$.
  • Delete one instance of $z$ from $E_{k}$.
  • Create $E_{k+1}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $E$ and adding 1 to each of them.
For the rest of this proof, let $\sum E_{k}$ denote the sum of the elements of $E_{k}$.
@@
If $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$, then the final number of bad edges in the graph is $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ for all $\varepsilon=\omega\left(n^{-1/2}\right)$.
@@
Let $T_{k}$ denote the number of edges that are incident to two vertices in $S_{k}$. Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum E_{k}$ denote $(\sum E_{k})-(\sum E_{k-1})$. For every move that we play on an element $z$ in $E_{k}$, the sum $\sum E_{k}$ grows by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges that are contained inside of $S_{k}$ grows by $z$. This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$. We have
$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum E_{k}$
$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum E_{k}$
$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum E_{k}$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum E_{m}-\sum E_{0}.$
This lemma requires that $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$.
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}m^{2}|S_{0}|\right)-\sum E_{0}$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}+\Theta(n)+\varepsilon T_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)+\Theta(n)+\varepsilon T_{0}$
$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+\Theta\left(\varepsilon^{-1}n\right)+T_{0}.$
Since $\Omega\left(\varepsilon^{-1}\right)$ bad edges are required to play the first uncertain best response, we must have $S_{0}=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon n^{2}\right)$ must also be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. If $\varepsilon=\omega\left(n^{-1/2}\right)$, then $O\left(\varepsilon n^{2}\right)$ will dominate $\Theta\left(\varepsilon^{-1}n\right)$.
@@
Any vertex in $G$ is either in $S_{m}$, or $L$. This means that the only bad edges that are not counted by $T_{m}$ must be incident to a vertex in $S_{m}$ and a vertex in $L$. The edges that are incident to a vertex in $S_{m}$ and $G\setminus S_{m}=L$ are entirely described by $E_{m}$. Thus, the number of bad edges that exist between a vertex in $S_{m}$ and a vertex in $L$ is given by $\sum E_{m}$, which is $O\left(\varepsilon^{2}m^{2}\right)$. This is dominated by $T_{m}=O\left(\varepsilon n^{2}\right)+T_{0}$, so the asymptotic growth rate of the number of bad edges in the graph is given solely by that of $T_{m}$.
@@
It remains to show that $T_{0}$ cannot grow faster than $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)\geq\Theta(\varepsilon n^{2})$. By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$. By Lemma 6, we start with $O\left(\varepsilon^{-2}\right)$ bad edges, so $T_{0}=O\left(\varepsilon^{-2}\right)$. This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega\left(n^{-1/2}\right)\geq\Theta\left(n^{-2/3}\right)$. ∎

### v2

@@ 4.1 Proof of Improved Upper Bound
With our construction, Proposition 6 allows us to contain the bad edges in the graph at any step inside one of $S_{0},S_{1},\dots,S_{m}$. We will now present a combinatorial abstraction that will allow us to focus on the effect of the first best response played by each vertex in $P$.
For every integer $k$ satisfying $0\leq k\leq m$, we can use $S_{k}$ to construct a multiset $A_{k}$ to encode information about certain bad edges in $S_{k}$. Specifically, each element of $A_{k}$ will describe the number of edges between a particular vertex in $V\setminus S_{k}$ and any vertex in $S_{k}$, and each vertex in $V\setminus S_{k}$ with at least one connection to a vertex in $S_{k}$ will be encoded exactly once in such a manner.
Example 8.
Suppose that there existed 4 edges between a vertex in $S_{0}$ and some vertex $x_{1}\not\in S_{0}$, and 3 edges between a vertex in $S_{0}$ and some different vertex $x_{2}\not\in S_{0}$. Also, suppose that there existed no other edges between a vertex in $S_{0}$ and a vertex in $V\setminus S_{0}$. Then, we would have
$A_{0}=\{4,3\}.$
$S_{0}$$x_{1}$$x_{2}$
Figure 1: Diagram for Example 8.
Consider the state of the graph right before $p_{k}$ plays its first best response. By Proposition 6, all edges between $p_{k}$ and a vertex in $V\setminus S_{k-1}$ must be good. If we let $z$ be the number of edges incident to $p_{k}$ and a vertex in $S_{k-1}$, then the degree of $p_{k}$ cannot be greater than $z+\lfloor(1+\varepsilon)z\rfloor$. By this logic, going from $A_{k-1}$ to $A_{k}$ consists of deleting the element of $A_{k-1}$ corresponding to the edges between a vertex in $S_{k-1}$ and $p_{k}$, which is $z$. Then, we will select $\lfloor(1+\varepsilon)z\rfloor$ elements of $A_{k-1}$ and increment them by $1$ (where we temporarily pad $A_{k-1}$ with zeroes to allow for the creation of new elements equal to 1).
Example 9.
Suppose that $A_{0}=\{3,2,2\}$, $\varepsilon=1/4$, and we wish to play a move on $z=4$. Then, $\lfloor(1+\varepsilon)z\rfloor=5$, and the possible multisets for $A_{1}$ are:
$\{3,3,1,1,1\},\quad\{3,2,1,1,1,1\},\;\;\text{and}\;\;\{2,2,1,1,1,1,1\}.$
The new multiset $A_{k}$ represents the edges between $p_{k}$ and a vertex in $V\setminus S_{k}$. Since any graph and sequence of best responses can be used to construct the multisets $(A_{k})_{0\leq k\leq m}$ in such a manner, it suffices to study the behavior of these moves.
@@
A regular move on a multiset $A_{k}$ with a positive integer $z\in A_{k}$ consists of the following steps:
  • Delete one instance of $z$ from $A_{k}$.
  • Pad $A_{k}$ with an sufficiently large (at least $z(1+\varepsilon)$) number of zeroes.
  • Create $A_{k+1}$ from $A_{k}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $A_{k}$ and adding 1 to each of them.
  • Remove all zeroes from $A_{k+1}$.
As a shorthand notation, let $\sum A_{k}$ denote the sum of the elements (with their multiplicities) of $A_{k}$, and let $\sum A_{k}^{2}$ denote the sum of squares of the elements (with their multiplicities) of $A_{k}$. The former will be a useful measure to bound because of Lemma 12, and the latter will assist us in obtaining such a bound.
Proposition 11.
$\sum A_{0}=O\left(n|S_{0}|\right)$.
Proof.
Since there are only $n$ vertices in the graph, every vertex in $S_{0}$ can have at most $n$ edges. Recall that $\sum A_{0}$ counts the number of edges between a vertex in $S_{0}$ and a vertex in $V\setminus S_{0}$. Proposition 11 follows. ∎
@@
If $\sum A_{m}=O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$, then the final number of bad edges in the graph is $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ if $\varepsilon=\omega(n^{-1/3})$.
@@
Let $T_{k}$ denote the number of edges between two vertices in $S_{k}$. Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum A_{k}$ denote $(\sum A_{k})-(\sum A_{k-1})$. For every regular move that we play on $A_{k-1}$ with an integer $z\in A_{k-1}$, the sum $\sum A_{k}$ grows from $\sum A_{k-1}$ by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges between two vertices in $S_{k}$ grows from $S_{k-1}$ by $z$. This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$. We have
$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum A_{k}$
$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum A_{k}$
$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum A_{k}$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum A_{m}-\sum A_{0}.$
Since $\sum A_{m}=O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$,
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}.$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}+\Theta(n)+\varepsilon T_{0}$
$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)-\varepsilon^{-1}\sum A_{0}+\Theta\left(\varepsilon^{-1}n\right)+T_{0}.$
By Proposition 11, $\varepsilon^{-1}\sum A_{0}$ must be $O(\varepsilon^{-1}n|S_{0}|)$. Since $\varepsilon=\omega(n^{-1/3})$, this is dominated by the $O(\varepsilon^{2}n^{2}|S_{0}|)$ term. Also, since we must have $\Omega(\varepsilon^{-1})$ starting bad edges, $S_{0}$ is $\Omega(\varepsilon^{-1})$ as well. This means that $O(\varepsilon n^{2})$ must also be $O(\varepsilon^{2}n^{2}|S_{0}|)$. Finally, $\varepsilon=\omega(n^{-1/3})$, so $O(\varepsilon n^{2})$ will dominate $\Theta(\varepsilon^{-1}n)$. We are left with the following terms on the RHS:
@@
We will now show that $T_{0}$ cannot grow faster than $O(\varepsilon^{2}n^{2}|S_{0}|)$. By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$. By Proposition 5, we start with $O(\varepsilon^{-2})$ bad edges, so $T_{0}=O(\varepsilon^{-2})$. This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega(n^{-1/3})$. Then, $O(\varepsilon n^{2})$ is $O(\varepsilon^{2}n^{2}|S_{0}|)$ because $|S_{0}|=\Omega(\varepsilon^{-1})$. Our equation finally becomes
$T_{m}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right).$
Any vertex in $V$ is either in $S_{m}$, or $Q$. This means that the only bad edges that are not counted by $T_{m}$ must be between $S_{m}$ and a vertex in $Q$, which are entirely described by the elements in $A_{m}$. Thus, the number of bad edges between a vertex in $S_{m}$ and a vertex in $Q$ is bounded by $\sum A_{m}$, which is $O(\varepsilon^{3}n^{2}|S_{0}|)$. The total number of bad edges is then bounded by
$T_{m}+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)=O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)=O\left(\varepsilon^{2}n^{2}|S_{0}|\right).$
Intuitively, playing a regular move on $\sum A_{k-1}$ with a large $z$ would result in a big increase in from $\sum A_{k-1}$ to $\sum A_{k}$. However, playing such a move might also create many small elements in $A_{k}$. Having the sum $\sum A_{k}$ spread out across many small elements will make increasing $\sum A_{k}$ in future moves more difficult. We consider $\sum A_{k}^{2}$ as a way to measure how ”spread out” the multiset $A_{k}$ is, and this particular quantity is preferable over other measures because the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ can be related to $\sum A_{k-1}$.
Our plan is to bound the behavior of $\sum A_{k}$ and $\sum A_{k}^{2}$ as $k$ increases. The multisets $A_{0}$ and $A_{m}$ will be subject to the following inequalities:
$\sum A_{m}^{2}$ $\geq 0$ (1)
$\sum A_{0}^{2}$ $\leq\left(\sum A_{0}\right)^{2}.$ (2)
Eq. 1 is true because the elements of $E_{m}$ must be nonnegative integers, and Eq. 2 follows from the Cauchy-Schwarz inequality. We will only consider these inequalities (and not those with $m$ and 0 swapped) because we only need to prevent too many regular moves from being played with large $z$, which will naturally cause $\sum A_{k}^{2}$ to become small as $k$ increases.
We will begin by investigating the behavior of the sum of squares. Let the current multiset be $A_{k-1}$. Then, pick some element $z\in A_{k-1}$, and define $\zeta=\lfloor z(1+\varepsilon)\rfloor$. We can play a regular move on $A_{k-1}$ with $z$, which will involve incrementing $\zeta$ the elements $a_{1},a_{2},\dots,a_{\zeta}\in A_{k-1}$ to $a_{1}+1,a_{2}+1,\dots,a_{\zeta}+1\in A_{k}$. It is easy to see that the effect of such a move on the increase from $\sum A_{k-1}$ to $\sum A_{k}$ is
$\sum A_{k}-\sum A_{k-1}$ $=-z+\sum_{i=1}^{\zeta}\big((a_{i}+1)-a_{i}\big)$
$\sum A_{k}-\sum A_{k-1}$ $=-z+\zeta.$ (3)
Next, we must have
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+\sum_{i=1}^{\zeta}\left((a_{i}+1)^{2}+a_{i}^{2}\right)$
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta.$ (4)
Eq. 3 and Eq. 4 together describe a single regular move. However, we can establish bounds upon these equations that are easier to analyze. Since $a_{1},a_{2},\dots,a_{\alpha}$ are all elements of $A_{k-1}$, their sum must be bounded by $\sum A_{k-1}$. By plugging this into our expression for the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$, we get
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta$
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum_{i=1}^{\zeta}a_{j}-z^{2}$
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum A_{k-1}-z^{2}.$ (5)
Proposition 13.
Taking the upper bound in Eq. 5 for the increase from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ never decreases $\sum A_{m}$.
Proof.
Notice that the changes described in the right hand sides of Eq. 3, Eq. 4, and Eq. 5 are all independent of the current value of $\sum A_{k-1}^{2}$. Furthermore, the only constraint on the final sum of squares $\sum A_{m}^{2}$ is a lower bound provided by Eq. 1. This means that always taking the upper bound for $\sum A_{k}$ from Eq. 5 will never decrease $\sum A_{m}$. ∎
Now equipped with Proposition 13, we actually no longer need that $z\in A_{k-1}$. In fact, we no longer even need to consider $A_{k-1}$ as a multiset anymore. For the rest of the proof, we will simply assume that $z$ is a positive real number. The next step is to find an upper bound for the change from $\sum A_{k-1}$ to $\sum A_{k}$.
$\sum A_{k}-\sum A_{k-1}$ $=\zeta-z$
$\sum A_{k}-\sum A_{k-1}$ $=\lfloor z(1+\varepsilon)\rfloor-z$
$\sum A_{k}-\sum A_{k-1}$ $\leq z(1+\varepsilon)-z+1$
$\sum A_{k}-\sum A_{k-1}$ $\leq\varepsilon z+1.$ (6)
Proposition 14.
Taking the upper bound in Eq. 6 for the increase from $\sum A_{k-1}$ to $\sum A_{k}$ never decreases $\sum A_{m}$.
Proof.
Suppose that we make a move using updates for the sum and sum of squares described by Eq. 3 and Eq. 5 respectively with some value for $z$. Then, we will show that there always exists another value $z^{\prime}$ with which we can make a move using Eq. 6 and Eq. 5 that results in the same value of $\sum A_{k}$ but a larger value of $\sum A_{k}^{2}$. By the same logic that was used in the proof of Proposition 13, this will never decrease $\sum A_{m}$.
Since the values of $\sum A_{k}$ must match, our selection of $z^{\prime}$ must satisfy
$\varepsilon z^{\prime}=\zeta-z=\lfloor z(1+\varepsilon)\rfloor-z.$
The RHS of this equation is clearly positive, so $z^{\prime}$ must be nonpositive as well. We can also bound the size of $z^{\prime}$ with
$\varepsilon z^{\prime}$ $=\lfloor z(1+\varepsilon)\rfloor-z$
$\varepsilon z^{\prime}$ $\leq z(1+\varepsilon)-z$
$z^{\prime}$ $\leq z.$
By (5), smaller values of $z$ yield larger values of $\sum A_{k}^{2}$. This is exactly what we wanted. ∎
With Proposition 13 and Proposition 14, we now have a strictly stronger set of actions on $\sum A_{k-1}$ and $\sum A_{k-1}^{2}$ that we can instead use with each of our moves.
Definition 15.
A strong move with $z$ on $(\sum A_{k-1},\sum A_{k-1}^{2})$ makes the following updates:
$\sum A_{k}$ $=\sum A_{k-1}+\varepsilon z+1$
$\sum A_{k}^{2}$ $=\sum A_{k-1}^{2}+2\sum A_{k-1}-z^{2}.$
For all integers $1\leq k\leq m$, let $z_{k}$ denote the value of $z$ used in the strong reversed move from $(\sum A_{k-1},\sum A_{k-1}^{2})$ to $(\sum A_{k},\sum A_{k}^{2})$. Then,
$\sum A_{k}$ $=\sum A_{0}+k+\varepsilon\sum_{i=1}^{k}z_{i}.$ (7)
The same process of expanding is more complicated for $\sum A_{m}^{2}$:
$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{i}\right)-\sum_{i=1}^{m}z_{i}^{2}.$
We can rewrite the $\sum A_{j}$ by plugging in Eq. 1 and Eq. 7.
$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{0}+i+\varepsilon\sum_{j=1}^{i}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
$0$ $\leq\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\varepsilon\sum_{j=1}^{m}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum_{i=1}^{m}z_{i}^{2}$ (8)
By the Cauchy-Schwarz Inequality,
$\left(\sum_{j=1}^{m}1\right)\left(\sum_{j=1}^{m}z_{j}^{2}\right)$ $\geq\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
$\sum_{j=1}^{m}z_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}.$
Plugging this result into Eq. 8 yields a quadratic polynomial in terms of $\sum_{i=1}^{m}z_{i}$ in the RHS.
$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
$0$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}-2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum A_{0}^{2}$
Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}z_{j}$ must be less than the larger root of this quadratic. This larger root is obtained through a tedious application of the quadratic formula:
$\sum_{j=1}^{m}z_{j}\leq\frac{2\varepsilon m+\sqrt{4\varepsilon^{2}m^{2}+4m^{-1}\sum A_{0}^{2}}}{2m^{-1}}$ (9)